The Super Discount store (open 24 hours a day, every day) sells 8-packs of paper towels, at the rate of approximately 420 packs per week. Because the towels are so bulky, the annual cost to carry them in inventory is estimated at $.50. The cost to place an order for more is $20 and it takes four days for an order to arrive.

a. Find the optimal order quantity.

b. What is the reorder point?

c. How often should an order be placed? 


1
Expert's answer
2022-01-28T12:48:01-0500

Weekly demand "=420"

Daily demand "=\frac{420}{7}=60"

Number of days for an order to arrive or lead time "=4"

Annual demand, "D=420 \times 52=21840"

Carrying cost, "H= \$0.50"

Ordering cost, "S=\$20"


a.)

The optimal order quantity, EOQ can be calculated as:


"EOQ=\sqrt{\frac{2DS}{H}}\\
=\sqrt{\frac{2\times 21840 \times 20}{0.50}}\\
\approx 1322"

b.)

Calculating the reorder point, "R":


"R= (\text{Average daily usage rate } \times \text{ Lead time})+\text{ Safety stock}\\
R=(60 \times 4)+0\\
R=240"

c.)

Frequency in which order should be placed:


"\text{Time}=\frac{365 \times EOQ}{D}\\
=\frac{365 \times 1322}{21840}\\
\approx 22 \text{ days}"